The no.of hydronium ions present in 10ml of water at 25 centigrade ?

Grade:12

1 Answers

reshma lr
10 years ago
pH=7 7=-log[H+] con of H+=10^-7 n=cv =10^-7*10*10^-3 =10^-9moles no of ions=6.o22*10^23*2*10^-9 =12.044*10^14 hope this helped you